On July 1, 2019, Wildhorse Co. purchased new equipment for $90,000. Its estimated useful life was 8 years with a $10,000 salvage value. On December 31, 2022, the company estimated that the equipment's remaining useful life was 10 years, with a revised salvage value of $5,000.


 Prepare the journal entry to record depreciation on December 31, 2019.

 Compute the revised annual depreciation on December 31, 2022. 

Compute the balance in Accumulated Depreciation Equipment for this equipment after depreciation expense has been recorded on December 31, 2022.
Answer #1

a) Journal entry

Account title and explanation Debit Credit
Depreciation expense (90000-10000/8)*6/12 5000
Accumulated depreciation-equipment 5000

b) Journal entry

Account title and explanation Debit Credit
Depreciation expense 10000
Accumulated depreciation-equipment 10000

c) Revised depreciation = (90000-25000- 5000)/10 = 6000

d) Journal entry

Account title and explanation Debit Credit
Depreciation expense 6000
Accumulated depreciation-equipment 6000

e) Accumulated depreciation- equipment = 25000+6000 = $31000